OpenStudy (anonymous):

A solid gold cube with edge of 2" is melted down and recast as a cone with a 3" radius base. Find the height of the cone.

OpenStudy (anonymous):

its simple, the volumes would be the same for cube its s^3 = 2^3 = 8 and for cone its (1/3)(pi)(r^2)(h) = 8 (pi)(3^2)(h)=24 pi(h)=24/9=8/3 h= (8/3)3.14 I hope its correct

OpenStudy (anonymous):

thnks it is

OpenStudy (anonymous):

welcome :D

OpenStudy (anonymous):

Find the surface area of a sphere having a radius of 4". HELP

OpenStudy (anonymous):

its just a formula???

OpenStudy (anonymous):

4(pi)(r^2)

OpenStudy (anonymous):

I got 64

OpenStudy (anonymous):

its 64(pi) right??

OpenStudy (anonymous):

yes